OpenStudy (jvillegas):

what is the recursive formula when given the explicit formula for the following geometric sequence? an= 12(33)^n-1

OpenStudy (loser66):

any idea?

OpenStudy (jvillegas):

nope /:

OpenStudy (loser66):

then, please calculate \(a_1\\a_2\\a_3\\a_4\)

OpenStudy (jvillegas):

is it an=33an-1?

OpenStudy (loser66):

yes, it is. not know how you get it, but it is correct

OpenStudy (zzr0ck3r):

\(a_n=12*33^{n-1}\) \(a_{n+1}=12*33^n\) If I am doing this right, \(a_{n+1}=33*a_n\) Does this make sense

OpenStudy (zzr0ck3r):

ignore me maybe...

OpenStudy (jvillegas):

LMAO. thank you both!

OpenStudy (zzr0ck3r):

oh, same thing...